find the three consecutive odd integers whose sum is 87
Answer by checkley77(12844)

LET X, X+2 & X+4 BE THE 3 ODD INTEGERS.
X+(X+2)+(X=4)=87
3X+6=87
3X=87-6
3X=81
X=81/3
X=27 ANSWER FOR THE SMALLER INTEGER.
27=2=29 FOR THE MIDDLE INTEGER.
27+4=31 FOR THE LARGEST INTEGER.
PROOF
27+29+31=87
87=87